OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4).

OpenStudy (anonymous):

f(1) = 1^2 + 1 = 2 and g(4) = 3(4) + 1 = 13 2f(1) + 3g(4) = 2(2) + 3(13) = 4 + 39 = 43

OpenStudy (anonymous):

It's a 'plug-and-chug' sort of problem. 2f(1) of x translates to \[2*f(x=1) = 2* [(1)^2 +1] = 2* (1 +1) = 4\] \[3g(4)=3*g(x=4) = 3*[3(4)+1] = 3*(12+1)=39\] Therefore \[2f(1) +3g(4) = 4 +39 = 43\]